OpenStudy (anonymous):

x/2 + x/4 = 5 i got 20, first i multiplied * 2 x + x/4 = 5 * 2 then i multiple by 4 x + x = 5*2*4 2x =40 then divide by 2 and got 20.. but when i checked my awnser, it was wrong where did i go wrong?

OpenStudy (anonymous):

You forgot to multiply x/4 by 2 in that first multiplication

OpenStudy (whpalmer4):

\[\frac{x}{2} + \frac{x}{4} = 5\]Multiply by 2: \[2*\frac{x}{2} + 2*\frac{x}{4} = 2*5\]\[x + \frac{x}{2} = 10\]Now multiply by 2 again and you should be in good shape...

OpenStudy (whpalmer4):

If you checked your answer by plugging your answer into the original problem and seeing if it worked (as opposed to entering it in your online homework system and getting it marked wrong), you get my stamp of approval for good working practices!

OpenStudy (anonymous):

i dont have a online homework system, it was on paper so after i multiply by 2 it would become 2x/8 = 10?

OpenStudy (anonymous):

Seems not, i'm confused how is 2 * x/4 = x/2 @whpalmer4

OpenStudy (whpalmer4):

If you have 2 quarters of something, isn't that the same as having 1 half of it? \[2*\frac{x}{4} = \frac{x}{4}+\frac{x}{4} = \frac{x+x}{4} = \frac{2x}{4} = \frac{2x}{2*2} = \frac{\cancel{2*}x}{\cancel{2*}2} = \frac{x}{2}\]